1. The product of the place values of two 2’s in 428721 is

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student
There are two 2’s in the given number.
the first 2 is in the tenth place,
Then, the product is = 2 × 10
= 20

The other 2 is in place value of ten thousand.

= 2 × 10000 = 20000

Thus, the product of place values = 20 × 20000 = 400000

Thanks